# On the uncertainty in predicting the stochastic gravitational wave background from compact binary coalescences

### v1

@@ II.1 Fiducial binary populations
For the BBH population, we adopt the best-fit mass distribution inferred from GWTC-3 based on the confident detections with an inverse false alarm rate (iFAR) greater than one year. Specifically, we use the preferred power-law plus peak (PLP) model described in the appendix of Ref. [11]. In this model, the primary mass distribution follows a power law with spectral index $a$ between $m_{\mathrm{min}}$ and $m_{\mathrm{max}}$ with an additional Gaussian peak with mean $m_{\mu}$ and width $m_{\sigma}$. The mixing fraction $\lambda_{p}$ determines the relative contribution of the Gaussian peak and the power law component. Additionally, the parameter $\delta_{m}$ governs the smoothing at the lower end of the mass spectrum. The mass ratio $q=m_{2}/m_{1}$ distribution follows a power law with spectral index $\beta_{q}$. The hyperparameters defining the fiducial PLP model are set to: a=3.5a=-3.5, $_{\mathrm{min}}=5.1\,M_{\odot}, mm$$_{\mathrm{max}}=88\,M_{\odot}$, $\lambda_{p}=0.038$, $m_{\mu}=34\,M_{\odot}$, $m_{\sigma}=4.6\,M_{\odot}$, $\delta_{m}=5.0\,M_{\odot}$ and $\beta_{q}=1.1$, based on the median values from the population inference presented in Ref. [11]. Given that the observed BBH population exhibits low effective inspiral spins [11, 62], we assume non-spinning BBHs in our fiducial model. Nonetheless, in Section III.2 we explore the impact of alternative spin distributions.
@@ III.1 BBH mass models
In Section II.1, we introduced the PLP mass distribution model, which provides the best fit to the observed BBH population in GWTC-3 [11]. For our fiducial model we adopted a single realization of this model, using the median values of the hyperparameters inferred from the population analysis. To accurately capture the uncertainties and correlations among these hyperparameters, we incorporate the full set of publicly available posterior samples obtained from the hierarchical Bayesian analysis of BBH mergers in GWTC-3. For each sample, we compute the corresponding GWB energy spectrum. The resulting ensemble of spectra, shown in Fig. 2, reflects the full uncertainty associated with the mass distribution. Additionally, we highlight the 90% credible band arising solely from the uncertainty in the mass distribution.
@@
Figure 3: Comparison of the 90% credible bands of the BBH background spectrum for different mass distribution models.
@@ III.2 BBH spin models
We investigate the impact of different spin models on the BBH background energy spectrum and compare them to the non-spinning fiducial models. The GWTC-3 population analysis considered two spin models [11]. The first, referred to as the default model, treats spin magnitudes and orientations independently. Spin magnitudes for both black holes are drawn from a Beta distribution, while tilt angles, defined as the angle between each black hole’s spin and the orbital angular momentum, are drawn from a mixture of an isotropic distribution and a preferentially aligned component, modeled by a truncated Gaussian distribution [77]. The second, known as the Gaussian spin model, describes the joint distribution of the effective inspiral spin parameter $\chi_{\mathrm{eff}}$ and the effective precession spin parameter $\chi_{\mathrm{p}}$ using a bivariate Gaussian with non-zero correlation [62]. Using posterior samples of the spin model hyperparameters from the population inference [11], we compute the GWB energy spectrum for each sample, modeling spins according to the distributions while keeping all other parameters fixed to those in the fiducial model. To account for spin precession, we employ the waveform model IMRPhenomXP [78]. The influence of spin distributions on the resulting background energy spectrum is notably smaller than that of mass distribution uncertainties. Therefore, we present only the relative difference with respect to the fiducial model in Fig. 4. At low frequencies, both spin models cause amplitude uncertainties of approximately 3%, while at higher frequencies, the background amplitude generally increases, more prominently for the Gaussian spin model. At 200 Hz, the enhancement can reach up to 10%.
@@ III.3 Impact of higher-order modes
So far we have computed the background energy spectra using only the dominant (2,2) mode of the waveform. Here, we assess the impact of including higher-order modes by employing the waveform model IMRPhenomXPHM. Since higher-order modes depend differently on the binary’s inclination angle than the (2,2) mode, it is necessary to explicity sample over inclination angles when computing $\Omega_{\mathrm{GW}}(f)$ via Eq. 2, rather than relying on an isotropic average. We first evaluate the fiducial model using IMRPhenomXPHM. As shown in Fig. 5, the inclusion of higher-order modes leads to a modest increase in the background amplitude, particularly at higher frequencies near the peak and fall-off of $\Omega_{\mathrm{GW}}(f)$. At lower frequencies, where most BBHs are in the inspiral phase, the effect is negligible. Next, we keep the PLP mass distribution but select a set of hyperparameters that favor asymmetric masses, specifically we choose one with a low value of $\beta_{q}$. Since higher order modes become more prominent in binaries with unequal masses [79, 80], we expect a stronger impact. Comparing the resulting amplitude increase compared to the same mass distribution evaluated without higher order modes, we find that the increase is greater than in the fiducial model, though still relatively small. Furthermore, we consider a model that retains the fiducial mass distribution but incorporates a spin model favoring large spin magnitudes and tilt angles near $90^{\circ}$. Such a configuration results in a population of strongly precessing binaries, where the GW signal is modulated and subject to mode mixing, thereby increasing the relevance of higher-order modes [81, 82]. Comparing the resulting spectrum to the one computed with only the dominant mode, we find a similar amplitude increase as for the fiducial case. This outcome is expected: while precession modulates the waveform, it has a limited effect on the total radiated GW energy.
@@ III.4 BNS mass models
Over a broad range of the relevant frequency band, BNS systems remain in the inspiral phase, during which the stochastic GWB follows the characteristic spectral dependence $\Omega(f)\sim f^{2/3}$. The amplitude of this background is directly influenced by the population-averaged chirp mass, where the chirp mass is the combination of component masses $\mathcal{M}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$. Since GW energy emitted by a compact binary scales approximately as $\mathcal{M}^{5/3}$, a higher average chirp mass results in a stronger background signal, particularly at frequencies below the merger regime. Moreover, the distribution of chirp masses affects the spectral shape near the peak of $\Omega(f)$. Lower chirp masses shift the peak to higher frequencies, while higher masses cause deviations from the $f^{2/3}$ scaling at lower frequencies. A broader chirp mass distribution tends to flatten the peak, whereas a narrow distribution produces a more sharply defined peak. This behavior is illustrated in Fig. 6, where we compare the energy density spectrum for the fiducial uniform mass distribution between 1 and 2.5 $M_{\odot}$ with that of a delta function at the corresponding average chirp mass of 1.73 $M_{\odot}$. The two models yield nearly identical amplitudes across most of the spectrum, with noticeable differences only near the peak. Additionally, we display how varying the minimum or maximum mass of the fiducial distribution alters the shape and amplitude of the resulting energy density spectrum.
@@ IV.3 Metallicity model
Figure 10: Impact of different modeling choices on the BBH merger rate and the resulting GWB energy density spectrum. Left panel: Effect of varying the time delay distribution. Shorter minimum delay times and steeper distributions lead to higher merger rates and enhanced background amplitudes. Middle panel: Influence of different SFR models. We compare the merger rates derived from the Madau-Fragos [65], Madau-Dickinson [64] and Vangioni [67] SFR models, shown both with (dashed lines) and without (solid lines) the fiducial metallictiy cut. While the choice of SFR has a minor effect, applying the metallicity cut substantially enhances the merger rate and the resulting GWB amplitude. Right panel: Effect of metallicity modeling. We compare the fiducial metallicity model using two different thresholds with the metallicity model from Ref. [98] evaluated for three thresholds, each with two values of metallicity dispersion. For reference we also show the case without any metallicity cut. Stricter metallicity thresholds and narrower dispersions shift the peak of the merger rate to higher redshifts and increase its amplitude, thereby enhancing the predicted GW background.
@@ V Combining uncertainties in modeling the astrophysical GWB
The resulting spectra are presented in Fig. 12, which illustrates both the individual realizations and the 90% credible band for the three classes of CBCs, alongside the fiducial models introduced earlier. In the bottom right panel of Fig. 12, we show again the 90% credible bands of all CBC backgrounds, together with the $2\sigma$ power-law integrated (PI) curves [106] from the stochastic cross-correlation search following the third observing run [18], as well as the projected sensitivity for one year of coincident observation at O5 sensitivity [107, 108] with upgraded Advanced LIGO detectors (A+) [109] and Virgo detector (V+) [110]. We note that while a detection of the combined background with O4 sensitivity is unlikely, we start probing the interesting regime, if the detectors reach the sensitivity projected for O5.
@@ Acknowledgements.
We would like to thank Irina Dvorkin for helpful comments. M.E. is supported by the Swiss National Science Foundation grant Sinergia 213497 and UZH Postdoc Grant, grant no. [FK-25-109]. We thank the Physik-Institut of the University of Zurich for providing computational resources.

### v2

# Uncertainty in predicting the stochastic gravitational wave background from compact binary coalescences

@@ II.1 Fiducial binary populations
For the BBH population, we adopt the best-fit mass distribution inferred from GWTC-3 based on the confident detections with an inverse false alarm rate (iFAR) greater than one year. While our analysis is based on GWTC-3 population inferences, we note that GWTC-4 [62, 63] has since been released, its potential impact on our results is discussed in Section VI. Specifically, we use the preferred power-law plus peak (PLP) model described in the appendix of Ref. [11]. In this model, the primary mass distribution follows a power law with spectral index $a$ between $m_{\mathrm{min}}$ and $m_{\mathrm{max}}$ with an additional Gaussian peak with mean $m_{\mu}$ and width $m_{\sigma}$. The mixing fraction $\lambda_{p}$ determines the relative contribution of the Gaussian peak and the power law component. Additionally, the parameter $\delta_{m}$ governs the smoothing at the lower end of the mass spectrum. The mass ratio $q=m_{2}/m_{1}$ distribution follows a power law with spectral index $\beta_{q}$. The hyperparameters defining the fiducial PLP model are set to: $a=3.5$, $_{\mathrm{min}}=5.1\,M_{\odot}, mm$$_{\mathrm{max}}=88\,M_{\odot}$, $\lambda_{p}=0.038$, $m_{\mu}=34\,M_{\odot}$, $m_{\sigma}=4.6\,M_{\odot}$, $\delta_{m}=5.0\,M_{\odot}$ and $\beta_{q}=1.1$, based on the median values from the population inference presented in Ref. [11]. Given that the observed BBH population exhibits low effective inspiral spins [11, 64], we assume non-spinning BBHs in our fiducial model. Nonetheless, in Section III.2 we explore the impact of alternative spin distributions.
@@ III.1 BBH mass models
In Section II.1, we introduced the PLP mass distribution model, which provides the best fit to the observed BBH population in GWTC-3 [11]. For our fiducial model we adopted a single realization of this model, using the median values of the hyperparameters inferred from the population analysis. To accurately capture the uncertainties and correlations among these hyperparameters, we incorporate the full set of publicly available posterior samples obtained from the hierarchical Bayesian analysis of BBH mergers in GWTC-3. For each sample, we compute the corresponding GWB energy spectrum. The resulting ensemble of spectra, shown in Fig. 2, reflects the full uncertainty associated with the mass distribution. Additionally, we highlight the 90% credible band arising solely from the uncertainty in the mass distribution. At lower frequencies, the background amplitude is proportional to the average chirp mass of the BBH population raised to the power of 5/3, where the chirp mass is the combination of component masses $\mathcal{M}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$. For the PLP mass distribution model, this average is primarily a function of the spectral index $a$ of the power law and $\lambda_{p}$, the fraction of binaries in the Gaussian peak. A steeper power law leads to more low mass binaries, while a smaller $\lambda_{p}$ reduces the contribution from the Gaussian peak near $34~M_{\odot}$. The exact location and width of the Gaussian peak have only a marginal effect. These two parameters also largely control the shape of the background spectrum near its maximum. While a steeper power law shifts the peak toward higher frequencies, a dominant Gaussian peak tends to broaden it toward lower frequencies.
@@
Figure 3: Comparison of the 90% credible bands of the BBH background spectrum for different mass distribution models. The solid line represents the median within each band. Note that the two GWTC-2 mass models (in blue and orange) show nearly complete overlap.
@@ III.2 BBH spin models
We investigate the impact of different spin models on the BBH background energy spectrum and compare them to the non-spinning fiducial models. The GWTC-3 population analysis considered two spin models [11]. The first, referred to as the default model, treats spin magnitudes and orientations independently. Spin magnitudes for both black holes are drawn from a Beta distribution, while tilt angles, defined as the angle between each black hole’s spin and the orbital angular momentum, are drawn from a mixture of an isotropic distribution and a preferentially aligned component, modeled by a truncated Gaussian distribution [79]. The second, known as the Gaussian spin model, describes the joint distribution of the effective inspiral spin parameter $\chi_{\mathrm{eff}}$ and the effective precession spin parameter $\chi_{\mathrm{p}}$ using a bivariate Gaussian with non-zero correlation [64]. Using posterior samples of the spin model hyperparameters from the population inference [11], we compute the GWB energy spectrum for each sample, modeling spins according to the distributions while keeping all other parameters fixed to those in the fiducial model. To account for spin precession, we employ the waveform model IMRPhenomXP [80].
The influence of spin distributions on the stochastic background energy spectrum is notably smaller than that of mass distribution uncertainties. Therefore, we present only the relative difference with respect to the fiducial model in Fig. 4. At low frequencies, both spin models cause amplitude uncertainties of approximately 3%, while at higher frequencies, the background amplitude generally increases, more prominently for the Gaussian spin model. Around 200 Hz, this enhancement can reach up to 10%. We also note that at low frequencies the spectrum amplitude tends to decrease when spin effects are included. This behavior arises because spins are preferentially aligned [11], which accelerates the inspiral and reduces the energy radiated at lower frequencies [81]. In contrast, aligned-spin binaries experience the orbital hang-up effect [82, 83], allowing tighter orbits just before merger and emit more GW energy at higher frequencies compared to non-spinning systems.
@@ III.3 Impact of higher-order modes
So far we have computed the background energy spectra using only the dominant (2,2) mode of the waveform. Here, we assess the impact of including higher-order modes by employing the waveform model IMRPhenomXPHM. Since higher-order modes depend differently on the binary’s inclination angle than the (2,2) mode, it is necessary to explicity sample over inclination angles when computing $\Omega_{\mathrm{GW}}(f)$ via Eq. 2, rather than relying on an isotropic average. We first evaluate the fiducial model using IMRPhenomXPHM. As shown in Fig. 5, the inclusion of higher-order modes leads to a modest increase in the background amplitude, particularly at higher frequencies near the peak and fall-off of $\Omega_{\mathrm{GW}}(f)$. At lower frequencies, where most BBHs are in the inspiral phase, the effect is negligible. Next, we keep the PLP mass distribution but select a set of hyperparameters that favor asymmetric masses, specifically we choose one with a low value of $\beta_{q}$. Since higher order modes become more prominent in binaries with unequal masses [84, 85], we expect a stronger impact. Comparing the resulting amplitude increase compared to the same mass distribution evaluated without higher order modes, we find that the increase is greater than in the fiducial model, though still relatively small. Furthermore, we consider a model that retains the fiducial mass distribution but incorporates a spin model favoring large spin magnitudes and tilt angles near $90^{\circ}$. Such a configuration results in a population of strongly precessing binaries, where the GW signal is modulated and subject to mode mixing, thereby increasing the relevance of higher-order modes [86, 87]. Comparing the resulting spectrum to the one computed with only the dominant mode, we find a marginal amplitude increase as compared to the fiducial case. This outcome is expected: while precession modulates the waveform, it has a limited effect on the total radiated GW energy.
@@ III.4 BNS mass models
Over a broad range of the relevant frequency band, BNS systems remain in the inspiral phase, during which the stochastic GWB follows the characteristic spectral dependence $\Omega(f)\sim f^{2/3}$. The amplitude of this background is directly influenced by the population-averaged chirp mass. Since GW energy emitted by a compact binary scales approximately as $\mathcal{M}^{5/3}$, a higher average chirp mass results in a stronger background signal, particularly at frequencies below the merger regime. Moreover, the distribution of chirp masses affects the spectral shape near the peak of $\Omega(f)$. Lower chirp masses shift the peak to higher frequencies, while higher masses cause deviations from the $f^{2/3}$ scaling at lower frequencies. A broader chirp mass distribution tends to flatten the peak, whereas a narrow distribution produces a more sharply defined peak. This behavior is illustrated in Fig. 6, where we compare the energy density spectrum for the fiducial uniform mass distribution between 1 and 2.5 $M_{\odot}$ with that of a delta function at the corresponding average chirp mass of 1.73 $M_{\odot}$. The two models yield nearly identical amplitudes across most of the spectrum, with noticeable differences only near the peak. Additionally, we display how varying the minimum or maximum mass of the fiducial distribution alters the shape and amplitude of the resulting energy density spectrum.
@@ IV.3 Metallicity model
Figure 10: Impact of different modeling choices on the BBH merger rate and the resulting GWB energy density spectrum. Left panel: Effect of varying the time delay distribution. Shorter minimum delay times and steeper distributions lead to higher merger rates and enhanced background amplitudes. Middle panel: Influence of different SFR models. We compare the merger rates derived from the Madau-Fragos [67], Madau-Dickinson [66] and Vangioni [69] SFR models, shown both with (dashed lines) and without (solid lines) the fiducial metallictiy cut. While the choice of SFR has a minor effect, applying the metallicity cut substantially enhances the merger rate and the resulting GWB amplitude. Right panel: Effect of metallicity modeling. We compare the fiducial metallicity model using two different thresholds with the metallicity model from Ref. [103] evaluated for two thresholds, each with two values of metallicity dispersion. For reference we also show the case without any metallicity cut. Stricter metallicity thresholds and narrower dispersions shift the peak of the merger rate to higher redshifts and increase its amplitude, thereby enhancing the predicted GW background.
@@ V Combining uncertainties in modeling the astrophysical GWB
The resulting spectra are presented in Fig. 12, which illustrates both the individual realizations and the 90% credible band for the three classes of CBCs, alongside the fiducial models introduced earlier. In the bottom right panel of Fig. 12, we show again the 90% credible bands of all CBC backgrounds, together with the $2\sigma$ power-law integrated (PI) curves [111] from the stochastic cross-correlation search following the third observing run [18], as well as the projected sensitivity for one year of coincident observation at O5 sensitivity [112, 113] with upgraded Advanced LIGO detectors (A+) [114] and Virgo detector (V+) [115]. The PI curve indicates the sensitivity to a power-law signal integrated over the entire frequency band at a given confidence level (here, $2\sigma$). If a signal lies above the PI curve, it would be detectable with at least $2\sigma$ confidence. We note that while a detection of the combined background with O4 sensitivity is unlikely, we start probing the interesting regime, if the detectors reach the sensitivity projected for O5. A detection or non-detection of a stochastic background carries important implications for population models. In the case of a non-detection, certain combinations of parameters governing the merger rate and source properties could be excluded. Nevertheless, because the background depends on numerous parameters and various modeling assumptions, isolating individual constraints remains difficult.
@@ VI Conclusion
Since the completion of this work, the GWTC-4 catalog [62] and updated population inferences [63] have been released. These updates primarily refine the underlying BBH mass model and tighten the local merger rate estimates. Their impact on the uncertainty in the predicted energy density spectra is expected to be minor, we anticipate slightly narrower uncertainty bands in Fig. 12 for BBH, BNS and NSBH. However, the dominant source of uncertainty remains the merger rate evolution. We note that the forecast in Fig. 7 of Ref. [116], which incorporates the updated population, reports even larger uncertainties for the BBH background. This difference arises because their approach estimates the low-redshift merger rate evolution directly from observed BBHs rather than using a time-delay and metallicity model, resulting in broader uncertainty compared to our modeling.
@@ Acknowledgements.
We would like to thank Irina Dvorkin for helpful comments. M.E. is supported by the Swiss National Science Foundation grant Sinergia 213497 and UZH Postdoc Grant, grant no. [FK-25-109]. We thank the Physik-Institut of the University of Zurich for providing computational resources. The data that support the findings of this article are openly available at [117].